Find the laurent series for (1-exp2z)/z


1
Expert's answer
2021-10-19T03:45:52-0400

"f(z)=\displaystyle{\sum_{n=0}^{\infin}}c_n(z-z_0)^n"


"c_n=\frac{f^{(n)}(z_0)}{n!}"

"z_0=0"


"1-e^{2z}=-2z+2z^2-4z^3\/3+2z^4\/3-4z^5\/15+..."


"\frac{1-e^{2z}}{z}=\frac{1}{z}(1-e^{2z})=-2+2z-4z^2/3+2z^3/3-4z^4/15+..."